OpenStudy (anonymous):

1/2 + (2/3 + 3/4) - (4/5 • 5/6) = ? what i have done is the operations in the parentheses and then done the math from left to right. always get the same, and wrong answer...

OpenStudy (aroub):

u have to find the lcd??

OpenStudy (anonymous):

yes it's 12

OpenStudy (anonymous):

i am doing something wrong with the order of operations or something

OpenStudy (aroub):

ok wait ill solve it

OpenStudy (anonymous):

thank you!

OpenStudy (aroub):

15/12?

OpenStudy (anonymous):

i made it 15/12 as well

OpenStudy (anonymous):

the question is multiple choice and these are the options A. 1/16 B. 17/27 C. 13/18 D. 7/9 E. 5/6 i go the same answer as you guys...

OpenStudy (anonymous):

haha, I got 15/12 as well, just so you guys know it's not just you.

OpenStudy (anonymous):

i did the same too haha

OpenStudy (aroub):

hahaha! mabye there is a wrong in the question or something

OpenStudy (anonymous):

true, the answer is C.... 13/18?? wtf haha

OpenStudy (anonymous):

answeris 31/12

OpenStudy (anonymous):

is that a multiplication in the last bracket?

OpenStudy (anonymous):

yes, the dot is multiplication

OpenStudy (anonymous):

okay I just did this...\[1/2+2/3+3/4-(20/30)=1/2+3/4=5/4\]

OpenStudy (anonymous):

i did exactly that... i swear it's the test that's wrong but it's an ACT test. they wouldn't make that mistake.

OpenStudy (anonymous):

no i mean 5/4

OpenStudy (anonymous):

yes 15/12 = 5/4

OpenStudy (anonymous):

*facepalm*

OpenStudy (anonymous):

that's still not one of the options... and the answer IS 13/18

OpenStudy (anonymous):

well it can't be, so they must've made a mistake. This is a practice I assume. Is there someone official you can ask? haha

OpenStudy (anonymous):

how did u get 13/18

OpenStudy (anonymous):

i didn't the book is telling me... it's a good thing i asked and know that i'm not retarded haha

OpenStudy (anonymous):

are you absolutely sure that you wrote out the problem correctly in your question?

OpenStudy (anonymous):

oh sorry 13/18 was the answer given

OpenStudy (anonymous):

i'll write it again one second

OpenStudy (anonymous):

1/2 + (2/3 + 3/4) - (4/5 x 5/6)

OpenStudy (anonymous):

it's the same.

OpenStudy (anonymous):

yeah that's the same. Hmmm...it must be a mistake. People make mistakes when writing up these books. It happens

OpenStudy (aroub):

there is something wrong in the question

OpenStudy (anonymous):

ok thanks everyone. ughh stupid book